OpenStudy (anonymous):

If the length of one base of a trapezoid is 5, the length of the other base is 2x + 7, and the length of the midsegment is 6x – 24, what is the value of x? Show equations and all work that leads to your answer.

OpenStudy (dumbcow):

midsegment is avg of 2 bases 6x-24 = (5+2x+7)/2 6x -24 = x+6 5x = 30 x = 6